# ProveRAG: Provenance-Driven Vulnerability Analysis with Automated Retrieval-Augmented LLMs

### v2

@@ 4 ProveRAG Methodology
ProveRAG works based on two main stages: The generation module and the evaluation module. Figure 1 shows the proposed architecture. Note that the Retr. LLM, Gen. LLM, and Eval. LLM are the same model. First, in the generation module, the Retr. LLM will load web content from external sources (starting from NVD) about a CVE vulnerability, and it will summarize the content with respect to the exploitation and mitigation information. Next, the Gen. LLM will look at the summaries and generates detailed exploitation and mitigation strategies for that specific CVE. Furthermore, in the evaluation module, the Eval. LLM will critique its own response by looking at the content from the official and external (NVD, CWE, and hyperlinks in NVD) websites to show evidence on whether its response were accurate, hallucinated, or omitted critical information. Finally, the LLM will provide provenance by showing similar pieces of information from both the response and the context while providing rationale to ensure trust. It is worth noting that we used the OpenAI’s gpt-4o-mini [25] model as the LLM for ProveRAG. This model’s training data knowledge is up to Oct. 2023, with 128,000 tokens for context window [25]. However, to show the proposed ProveRAG framework can be generalized across other LLMs, we compare the performance of gpt-4o-mini with Llama-3.1-8B [17] in Sec. 5.6. The Llama-3.1-8B model is selected as it has the same context window of 128,000 tokens, and knowledge cutoff until Dec. 2023 (no exposure to CVEs from 2024).
4.1 System Design
4.1.1 Generation Module:
Generation Prompt You are a cybersecurity expert. Consider the Relevant Information provided below and answer the Query. Relevant Information: [content from NVD/CWE/Relevant Refs.] Query: [CVE-xxx-xxx] Given the specified CVE-ID, please provide detailed answers to the following questions: 1. How can an attacker exploit this vulnerability? Provide a step-by-step description. 2. What are the recommended mitigation strategies for this vulnerability?
The Relevant Information is gathered from the summarizing retrieval technique (to answer RQ1), and we will show in Sec. 5.2.1, that it is the better retrieval technique in vulnerability analysis. The Gen. LLM will look at the retrieved information for a particular CVE and will provide the response. In the prompt-only experiment, the Relevant Information part is removed as we directly query the LLM about a specific CVE. Note that we put the temperature hyperparameter (which assigns the degree of randomness in LLM’s output) to zero as we aim to get the most deterministic responses from the LLM.
@@ 4.1.1 Generation Module:
We test on two retrieval techniques (to address RQ2). Figure 2 shows how each of these techniques is used. For the chunking technique, the LLM will use top-10 most similar chunks of 15,000 characters from the resources. Second, for the summarizing technique, the Retr. LLM goes through the NVD, then to the CWE and hyperlinks to summarize the content for the exploitation and mitigation questions The following shows the prompt used in Retr. LLM to assess each URL’s relevancy, and, if relevant, the LLM will summarize the URL content specifically to address either the Exploitation or the Mitigation question.
Relevancy Prompts Exploitation Prompt You are a cybersecurity expert. Your task is to analyze the provided URL content for the [CVE-xxx-xxx] and provide a detailed summary. Content: content Please follow the steps below: Step 1: Assess Relevancy - Does the content provide relevant information to describe how this [CVE-xxx-xxx] can be exploited? - Answer: [Yes/No] Step 2: Summarize Relevant Information - If the answer is "Yes" in Step 1: - Summarize the content with step-by-step description to exploit this vulnerability. - If the answer is "No" in Step 1: - Summary: NONE Mitigation Prompt You are a cybersecurity expert. Your task is to analyze the provided URL content for the [CVE-xxx-xxx] and provide a detailed summary. Content: content Please follow the steps below: Step 1: Assess Relevancy - Does the content provide relevant information to describe the potential mitigation strategies for this vulnerability? - Answer: [Yes/No] Step 2: Summarize Relevant Information - If the answer is "Yes" in Step 1: - Summarize the content to describe the potential mitigation strategies for this CVE. - If the answer is "No" in Step 1: - Summary: NONE
In the chunking technique, we split the content of all the URLs into smaller chunks and embed these chunks as vectors, and these embeddings are then indexed to facilitate efficient retrieval. We further show the performance for different models in Sec. 5.6. During runtime, the user’s query is first vectorized, and a similarity search is conducted against the indexed chunks. The top-10 results are retrieved and fed as context into the Gen. LLM’s prompt, which allows the model to answer the user’s query with the most relevant content. However, this approach requires a very large context window and can include redundant information that might mislead the LLM.
@@
4.1.2 Evaluation Module:
Evaluation Prompt For the [exploitation/mitigation] information of [CVE-xxx-xxx]: Please evaluate the response based on the provided evidence and select the appropriate value for the defined attributes: ————————- response: [LLM response] ————————- evidence: [content from NVD + CWE + Relevant Refs.)] ————————-
The LLM is tasked with providing responses for three attributes: value, rationale, and provenance. The following are these attributes for the Eval. LLM:
Evaluation Module Value prompt - Description: This attribute represents the accuracy of the response based on the provided context about the CVE-ID. - Allowed Values: - ‘TP’ (True Positive): The response fully and accurately reflects the information in the context. - ‘FP’ (False Positive): The response contains inaccurate or incorrect information that are not supported by the context. - ‘FN’ (False Negative): The response omits information that is present in the context. - Requirements: - Must strictly be one of the allowed values: ‘TP’, ‘FP’, or ‘FN’. - The selection must be based on an objective comparison between the response and the context. - Must follow the guidelines for Provenance. Rationale prompt - Description: This attribute provides the reasoning behind the selected "value" for the evaluation of the response.
Evaluation Module Provenance prompt - Detailed Instructions for Provenance based on the selected "value": - If "value" is ‘TP’: 1. Carefully compare the response with the context. 2. Identify and extract the key segments from both the response and the context that align perfectly. 3. Format the provenance as follows: - ‘response: "[Extracted segment from the response that matches the context]".’ - ‘context: "[Corresponding segment(s) from the context]".’ - If "value" is ‘FP’: 1. Review the response thoroughly to identify parts that are incorrect or not present in the context. 2. Format the provenance as follows: - ‘response: "[Incorrect or unsupported segment from the response]".’ - ‘context: "[The closest matching part from the context, or indicate ‘No corresponding information in context’ if none exists]".’ - If "value" is ‘FN’: 1. Examine the context to find essential information that is present but omitted in the response. 2. Format the provenance as follows: - ‘response: "[The entire response]".’ - ‘context: "[Relevant segment from the context that should have been matched]".’ Guidelines: - In all cases, ensure the provenance provided is clear, concise, and directly supports the selected "value". - The provenance should be an exact match or a direct comparison between the response and the context, following the formatting rules strictly.
For these attributes, we ask the LLM to provide a value: True Positive (TP) - indicating accurate prediction, False Positive (FP) - indicating the detection of hallucination, or False Negative (FN) - indicating failed retrieval given the response and the evidence (from NVD+CWE+hyperlinks). Next, the LLM is asked to provide a rationale for the selected value, and finally the provenance attribute where the LLM is tasked to show evidence for its response.
To answer RQ3, for each exploitation and mitigation response separately, we retrieve the top-10 chunks from the NVD, CWE, and hyperlinks with respect to the CVE-ID as evidence (see Figure 2’s Evaluation Query), and we compare them with the response using the Eval. LLM. It is worth noting that we do not integrate the use of memory when querying the LLM. Instead, all prompts are submitted in a single batch, without any retention of context between queries.
5 Evaluation of ProveRAG
Table 1: Comparison between Summarizing and Chunking techniques for retrieving the content for CVEs (Provenance Features: TP: accurate prediction, FP: detected hallucination, FN: failed retrieval)
@@ 5 Evaluation of ProveRAG
To answer each research question, we discuss the curation of the dataset for (RQ1) and provide results for different RAG techniques (RQ1 and RQ2). Furthermore, we support the efficacy of provenance by using evaluation metrics (RQ3), show how each URL attributed to the Gen. LLM’s response (RQ3), and finally we evaluate ProveRAG on another LLM (Llama-3.1) and we will show some specific examples of ProveRAG’s responses.
5.1 Dataset & Experiment Design
@@
To design the experiments for addressing RQ1, we curated a dataset of critical CVEs reported in 2024 (up until July 25). We focused on CVEs with a Common Vulnerability Scoring System (CVSS) score of 9.0 or higher, which indicates critical severity. From this dataset, we further refined our selection by including only those CVEs that had a direct mapping to the CWE database (as CWE contains information about potential mitigation strategies). Moreover, we leveraged the hyperlinks available on the NVD website, which provide additional relevant information for each CVE. These hyperlinks were gathered to further enrich the dataset used in the ProveRAG framework, to ensure that we are cross-referencing all available sources of information related to each CVE. In total, the curated dataset comprised of 482 CVEs, each mapped to corresponding CWEs and supplemented with relevant NVD resources. This dataset forms the basis for our temporal alignment and provenance analysis in the ProveRAG framework (as gpt-4o-mini model has training knowledge up to Oct. 23, 2023). This enables a thorough assessment of the LLM’s ability to generate accurate, up-to-date, and contextually relevant security recommendations.
5.2 Effectiveness of ProveRAG
5.2.1 Chunking vs. Summarizing Techniques
@@ 5.2 Effectiveness of ProveRAG
To answer RQ2, for the evaluation of retrieval techniques, we compare multiple chunking techniques by increasingly augmenting the Retr. LLM with more external resources vs. the summarizing technique (including all the relevant resources) to assess their capabilities in providing exploitation and mitigation strategies. To do a fair comparison between the chunking vs. summarizing technique, we only used NVD and CWE as evidence sources of information. Table 1 shows the overall results for these two techniques.
From the results, first we can see that for the prompt-only experiment, the performance is poor, for both exploitation and mitigation results. This shows the big limitation of LLMs’ knowledge with new, unseen data, as the entire dataset consists of 2024 vulnerabilities. This goes back to RQ1 and how we can deal with this problem. By integrating the retrieval techniques, we can see that the temporal alignment of the LLM is improved by the inclusion of the curated dataset, indicating that such an approach may alleviate limitations related to temporal alignment for cybersecurity applications.
For the exploitation results, all evaluated methodologies exhibited high accuracy, especially for NVD-Only which was limited to the information about the CVE’s exploitation on the NVD website. Another observation is that adding more data slightly decreases the chunking technique’s performance. In Contrast, for the more complex queries (i.e., mitigation information), significant variations are observed. The NVD-Only configuration achieves around 29% accuracy in providing mitigation strategies, which makes sense given that the NVD website lacks a dedicated mitigation section; yet, it does offer some recommendations for certain CVEs. When the CWE is added into the generation of results, we can see a big increase in TPs (from 139 to 262). However, adding additional URLs into the generation mix appears to diminish performance by about 7%, which suggests that an excess of links might degrade the quality of the retrieved data segments with the chunking technique. This indicates that providing excessive context can degrade performance when chunking is used.
On the other hand, the summarizing technique significantly outperforms the chunking technique when provided with the same set of links for data generation, showing an increase from 227 to 379 in TPs — a rise of over 30%. Therefore, these findings strongly support the conclusion that summarization is a better and more efficient technique. As a result, it has been incorporated into ProveRAG to help align the system to vulnerabilities discovered after training cut-off. Furthermore, this technique alleviates the context window problem (RQ2) for vulnerability analysis, as only the essential information (i.e., summaries) are put into the Gen. LLM’s prompt.
5.2.2 ProveRAG Results
To fully integrate ProveRAG and assess whether there are pieces of evidence present in the hyperlinks, we compare the Gen. LLM’s response with all NVD, CWE, and hyperlinks. Table 2 shows the results of ProveRAG (The ProveRAG-Aqua is discussed in Section 5.5 and and the corresponding improvement in ProveRAG performance). From the results, we observe a minor reduction in the performance regarding exploitation detection, paired with a marginal improvement in identifying mitigation measures. This suggests that while the relevant references do not completely enhance the utility of the data, they nevertheless ensure comprehensive coverage of pertinent information related to specific CVEs. Section 5.7 shows the Example of CVE-2024-0244, which illustrates the benefits of including hyperlinks in enhancing the understanding of mitigation strategies for a particular CVE. This is a solution for RQ2, as there could be as many links to be processed and summarized to put in the context window of LLMs.
Table 2: ProveRAG results for the 482 CVEs
@@ 5.2.2 ProveRAG Results
It is important to note that based on manual analysis of TP results, ProveRAG acted as a conservative analyst, only providing responses that had a high degree of confidence and explainability from the data sources provided to the model. However, there are some observed inaccuracies in the detection of FP and FN responses, which contribute to the slight decrease in the exploitation performance values (Table 2). Next, we show evidence on the quality of the provenance responses.
5.3 Evaluation Metrics for Provenance
To answer RQ3, and to evaluate provenance quality, we made use of the following metrics:
5.3.1 Attribution:
To assess the attribution of each URL summary (i.e., NVD, CWE, and hyperlinks) and how much each contributed to the generation of the results, we made use of the Captum library [18]. This library offers implementations of various perturbation- and gradient-based attribution algorithms for PyTorch models. Attribution methods allow for model explainability by assigning scores to each input feature, representing its impact on the model’s output. In our experiment, we compute attribution scores using feature ablation, a perturbation-based method that evaluates the model on altered inputs. This method replaces each feature with its corresponding value from the baseline feature vector to assess its influence. Equation 1 assesses the contribution of different features to the final result.
$\Delta y=y(x)-y(x_{\setminus F})$ (1)
Here, $y(x)$ represents the model’s output given the complete original input $x$. The term $x_{\setminus F}$ denotes the modified input where the feature $F$ (or a group of features) has been ablated—specifically, replaced with a baseline or reference value. The difference $\Delta y$ quantifies the change in the model’s output attributable to the ablation of the feature $F$.
5.3.2 Rouge-L and Cosine Similarity:
@@
To evaluate the quality of the Provenance results, we incorporated Rouge-L and Cosine Similarity (based on embeddings) metrics into our analysis. Rouge-L focuses on the longest common subsequence, which measures the sequence similarity between the generated responses and the reference texts. On the other hand, Cosine Similarity leverages a Sentence Transformer model, specifically the multi-qa-mpnet-base-dot-v1, which is optimized for semantic search [30]. This model utilizes a 512-context window to compare the semantic congruence between the response and the evidence part of the provenance, ensuring that the evaluation captures not just the textual, but also the contextual and semantic accuracy of the information provided.
5.4 Provenance Quality
@@
To answer RQ3, we applied Rouge-L and Cosine Similarity to measure the quality of both the provided responses and their supporting provenance. These metrics help us assess the alignment between the system’s outputs and verifiable evidence when evaluating TPs, FPs, and FNs. Figure 3 illustrates these values. As anticipated, TPs consistently achieve the highest scores across both metrics, indicating greatest overlap between the generated response and the supporting evidence, which underscores the effectiveness of ProveRAG. Note that the scores of FNs is close to that of TPs. This is expected because the FNs often just omitted some details in the response when comparing it with the source context. There is no value for FN in the exploitation part, as the LLM made no FN predictions in that category. Detailed examples are provided in Section 5.7 to contextualize these values.
@@
Furthermore, to analyze the effect of summary for each URL to the LLM’s response, we made use of the Captum library [18]. We calculate attribution by using the summaries from each relevant website as context for the generation prompt. If a website is deemed irrelevant, its attribution score is set to zero, as it does not provide usable information for the LLM. To facilitate comparison across CVEs, the attribution scores are normalized to unit norm. Figure 4 shows these attribution values. From the results, it is evident that for exploitation, the majority of information in the summaries originates from NVD. For mitigation summaries, most of CWE’s attention scores are higher than 0.9, whereas the values for NVD and the hyperlinks stay mostly unchanged. This indicates that the Gen. LLM put the most attention to the NVD summary for generating the exploitation response, and to the CWE summary for generating the mitigation strategies.
Figure 4: ProveRAG Attribution Results
5.5 Improvement w/ Structured Context
In Sec. 5.2.2, we demonstrated that ProveRAG achieves 95% and 80% TPs (with NVD, CWE, and all hyperlinks in the NVD website). Therefore, we raise the question on whether this can be further improved. Note that CWE is addressing mitigation at the weakness level, not the vulnerability level. Through our research, we found Aqua Vulnerability Database [29], which gives a highly structured and detailed context for each CVE’s exploitation steps and mitigation strategies. Thus, we experimented ProveRAG by purposefully including Aqua in the retrieval summarizing process, and into the evaluation module for provenance. The results are shown in the bottom row of Table 2. The results demonstrate the significant increase in the exploitation and mitigation detection of ProveRAG-Aqua, with 99% and 97% for TPs, respectively. This shows that the Aqua database plays a pivotal role in enhancing ProveRAG’s performance by providing more detailed and structured context. These results are significantly better than the baseline experiments, where prompt-only achieved just 6% accuracy for mitigation strategies, and the retrieval chunking technique reached about 47%.
Table 3: Critical CVEs Relevancy with respect to the URLs

### v3

@@ 2. What are the recommended mitigation strategies for this vulnerability?
We test on two retrieval techniques (to address RQ2). Figure 2 shows how each of these techniques is used. For the chunking technique, the LLM will use the top 10 most similar chunks of 15,000 characters from the resources to the query. In this technique, we split the content of all the URLs into smaller chunks and embed these chunks as vectors, and these embeddings are then indexed to facilitate efficient retrieval. Second, for the summarizing technique, the Retr. LLM iteratively retrieves information from the NVD, the corresponding CWE entry, and any additional referenced resources to check each source’s relevancy for the exploitation and mitigation questions. If relevant, it will summarize the content and passes it to the Gen. LLM. In AppendixB, we provide the engineered prompts for the summarizing retrieval technique.
We further show the performance for different models in Sec. V-E. During runtime, the user’s query is first vectorized, and a similarity search is conducted against the indexed chunks. The top-10 results are retrieved and fed as context into the Gen. LLM’s prompt, which allows the model to answer the user’s query with the most relevant content. However, this approach requires a very large context window and can include redundant information that might mislead the LLM.
@@
TABLE I: Comparison between Summarizing and Chunking techniques for retrieving the content for CVEs (Provenance Features: TP: accurate prediction, FP: detected hallucination, FN: failed retrieval)
@@
IV-A2 Evaluation Module:
{mybox}
[frametitle=Evaluation Prompt]
For the [exploitation/mitigation] information of [CVE-xxx-xxx], please evaluate the response based on the provided evidence and select the appropriate value for the defined attributes:
response: [LLM response]
evidence: [content from NVD + CWE + Relevant Refs.)]
As can be seen, the LLM is tasked with providing responses for three attributes, and those attributes are: value, rationale, and provenance. The complete prompts for these attributes are provided in the AppendixA. For these attributes, we ask the LLM to provide a value: True Positive (TP) - indicating accurate prediction, False Positive (FP) - indicating the detection of hallucination, or False Negative (FN) - indicating failed retrieval given the response and the evidence (from NVD+CWE+hyperlinks). Next, the LLM is asked to provide a rationale for the selected value, and finally the provenance attribute where the LLM is tasked to show evidence for its response. Note that for provenance, the document chunking technique is utilized as we need the actual information from reputable sources (not the LLM-generated summaries).
In addition, we ensure that the LLM for ProveRAG do not miss any essential information when finding evidence by utilizing almost the entire context window (128,000 tokens) for provenance. Therefore, for each exploitation and mitigation response separately, we retrieve the top-10 chunks from the NVD, CWE, and hyperlinks with respect to the CVE-ID as evidence, and we compare them with the response using the Eval. LLM. All prompts are submitted in a single batch, without any retention of memory between queries.
V Evaluation of ProveRAG
@@
To answer each research question, we discuss the curation of the dataset for (RQ1) and provide results for different RAG techniques (RQ1 and RQ2). Furthermore, we support the efficacy of provenance by using evaluation metrics (RQ3), and we evaluate ProveRAG on another LLM to assess generalizability. Finally, we will show specific examples to demonstrate ProveRAG’s effectiveness.
V-A Dataset & Experiment Design
@@
To address RQ1, we curated a dataset of critical CVEs reported in 2024 (up until July 25). We focused on CVEs with a Common Vulnerability Scoring System (CVSS) score of 9.0 or higher, which indicates critical severity. From this dataset, we further refined our selection by including only those CVEs that had a direct mapping to the CWE database (as CWE contains information about potential mitigation strategies). Note that CWE is addressing mitigation at the weakness level, not the vulnerability level.
Moreover, we leveraged the hyperlinks available on the NVD website for each CVE, which provide additional relevant information for each CVE. These hyperlinks were gathered to further enrich the dataset used in the ProveRAG framework, to ensure that we are cross-referencing all available sources of information related to each CVE. In total, the curated dataset comprised of 482 CVEs, each mapped to its corresponding CWEs and supplemented with relevant NVD resources. This dataset forms the basis for our temporal alignment and provenance analysis for ProveRAG since the gpt-4o-mini and llama-3.1-8B models have a knowledge cutoff in September and December 2023, respectively. This allows for a thorough assessment of the LLM’s ability to generate accurate, up-to-date, and contextually relevant security recommendations.
V-B Effectiveness of ProveRAG
V-B1 Chunking vs. Summarizing Retrievals
@@ V-B Effectiveness of ProveRAG
To answer RQ2, for the evaluation of retrieval techniques, we compare multiple chunking techniques by increasingly augmenting the Retr. LLM with more external resources versus the summarizing technique (including all relevant resources) to assess their abilities to provide accurate exploitation and mitigation strategies. To make a fair comparison between the chunking vs. summarizing technique, we only used NVD and CWE as evidence sources of information. Table I shows the overall results for these two techniques.
The prompt-only baseline exhibited poor performance for both exploitation and mitigation, demonstrating the limitations of LLMs with unseen vulnerabilities (as highlighted in RQ1). By integrating the retrieval techniques, we can see that the temporal alignment of the LLM is improved, indicating that such an approach alleviates temporal limitations for cybersecurity applications. For exploitation, all RAG configurations achieved high accuracy, though performance slightly declined when additional data were introduced with the chunking approach. Mitigation queries, however, revealed greater variation: the NVD-only setting achieved 29% accuracy, consistent with its limited mitigation content. Augmenting with CWE substantially increased #TPs (139→262), while adding further URLs reduced performance (7%) due to lower segment quality.
In contrast, the summarizing technique consistently outperformed chunking when provided with the same evidence set. By synthesizing key information across sources into concise summaries, it raised TPs from 227 to 379, over 30% improvement. These findings support the adoption of summarization over chunking in vulnerability analysis. Therefore, ProveRAG incorporates summarization as the default retrieval strategy for response generation. Importantly, this approach also alleviates the context-window constraint (RQ2), since only essential distilled knowledge, not raw segments, is passed to the Gen. LLM. Next, we provide the ProveRAG results by incorporating the entire proposed pipeline.
TABLE II: ProveRAG results across all evidence sources for the 482 CVEs
@@
V-B2 ProveRAG Results
To fully integrate ProveRAG and assess whether there are pieces of evidence present in the hyperlinks, we compare the Gen. LLM’s response with all NVD, CWE, and hyperlinks. Table II shows the results of ProveRAG. Note that the ProveRAG-Aqua is discussed in Section V-D and we discuss how more structured content can improve ProveRAG’s performance. From the results, we observe a minor reduction in the performance regarding exploitation detection, paired with a marginal improvement in identifying mitigation measures (compared with Table II). This suggests that while the relevant references do not completely enhance the utility of the data, they nevertheless ensure comprehensive coverage of pertinent information related to specific CVEs. Section V-F shows the example of CVE-2024-0244, which illustrates the benefits of including hyperlinks in enhancing the understanding of mitigation strategies for a particular CVE. This is a solution for RQ2, as there could be as many links to be processed and summarized to put in the context window of LLMs.
@@ V-B2 ProveRAG Results
It is important to note that based on our manual analysis of the TP results, ProveRAG acted as a conservative analyst, only providing responses that had a high degree of confidence and explainability from the data sources provided to the model. However, there are some observed inaccuracies in the detection of FP and FN responses. These inaccuracies often occur when the LLM is given lengthy context with many possible exploitation or mitigation details. In such cases, the model compares its response against all available information, leading it to flag deviations as FP or FN even when the response is largely correct. More refined, adaptive prompting that directs the model to focus on the most essential evidence can help reduce these misclassifications. Next, we show evidence on the quality of the provenance responses.
V-C Provenance Quality
@@
To answer RQ3, and to evaluate provenance quality, we made use of the Rouge-L and Cosine Similarity (based on embeddings) metrics into our analysis. Rouge-L focuses on the longest common subsequence, which measures the sequence similarity between the generated responses and the reference texts. On the other hand, Cosine Similarity leverages a Sentence Transformer model, specifically the multi-qa-mpnet-base-dot-v1, which is optimized for semantic search [35]. This model utilizes a 512-context window to compare the semantic congruence between the response and the evidence part of the provenance, ensuring that the evaluation captures not just the textual, but also the contextual and semantic accuracy of the information provided.
@@
These metrics help us assess the alignment between the system’s outputs and verifiable evidence when evaluating TPs, FPs, and FNs. Figure 3 illustrates these values. As anticipated, TPs consistently achieve the highest scores across both metrics by having the most overlap between the generated response and the supporting evidence, which underscores the effectiveness of ProveRAG. Interestingly, we observe that the scores of FNs are close to those of TPs. This is expected because the FNs often just omitted some details in the response when comparing it with the source evidence. Note that there is no value for FN in the exploitation part, as the LLM made no FN predictions in that category. Detailed examples are provided in Section V-F to contextualize these values.
@@
V-D Improvement w/ Structured Context
In Sec. V-B2, we demonstrated that ProveRAG achieves 95% and 80% TPs (with NVD, CWE, and all hyperlinks in the NVD website). Therefore, we raise the question on whether this can be further improved. Note that CWE is addressing mitigation at the weakness level, not the vulnerability level. Through our research, we found Aqua Vulnerability Database [20], which gives a highly structured and detailed context for each CVE’s exploitation steps and mitigation strategies. Thus, we experimented ProveRAG by purposefully including Aqua in the retrieval summarizing process, and into the evaluation module for provenance. The results are shown in the bottom row of Table II. The results demonstrate the significant increase in the exploitation and mitigation detection of ProveRAG-Aqua, with 99% and 97% for TPs, respectively. This shows that the Aqua database plays a pivotal role in enhancing ProveRAG’s performance by providing more detailed and structured context. These results are significantly better than the baseline experiments, where prompt-only achieved just 6% accuracy for mitigation strategies, and the retrieval chunking technique reached about 47%.
TABLE III: Critical CVEs Relevancy with respect to the URLs